# Can Authorship Attribution Models Distinguish Speakers in Speech Transcripts?

### v3

@@ 4.1 Experiment 1: Baselines
Table 2: Bootstrapped test performance (AUC) across all out-of-the-box (o) models for the BBN, LDC, and NormLDC encodings across all difficulty levels. Best performance for each encoding and difficulty level is bolded and second best, underlined, the differences of which are all statistically significant ($p<0.001$) except for ties. The largest standard error is 0.0004.
@@
For the TF-IDF baseline, the vectorizer was fit to the Reuters-21578 corpus, which contains 10,788 news documents and totals 1.3 million words (Lewis, 1997), [^10] We also tried fitting the vectorizer to the training set transcripts but found similar results so fit to Reuters to focus on out-of-the-box performance of written text models. the document-term matrix obtained for each transcript in the test set trials, and the cosine similarity calculated between each matrix in each trial. For the out-of-the-box PAN baseline, PANgramso was trained on the most recent openly available PAN authorship verification dataset, fanfiction stories (Bevendorff et al., 2020), and evaluated on the test set trials using cosine similarity.
@@
AUC score test set results are in Table 2 and EER results are in Table 8 in Appendix A. Focusing on the leftmost column (BBN ‘text-like’ encoding, ‘base’ difficulty), there is a clear ranking in performance: PANgramso achieves the highest performance followed by LUARo, SBERTo, and CISRo, with all scores well above chance (AUC = $0.5$). AdHominemo and TF-IDFo perform worst but still above chance. These initial results—considering only the ‘base’ setting for now—suggest at least some transfer from the text domain to the speech domain. However, we revisit this idea in Section 4.3 when discussing results on topic-controlled datasets.
@@ 4.3 Experiment 3: Topic
To test the effect of topic, after running Experiment 1 on the ‘base’ dataset, we ran it on the ‘hard’ and ‘harder’ datasets. Since performance of both LUARo (and LUARnormo) was highest on the NormLDC encoding previously, we included it here as well. The full AUC results across all settings are shown in Table 2 and the corresponding EER results are in Table 8.
@@ Appendix A Appendix
Table 8: Bootstrapped test performance (EER) across all out-of-the-box (o) models and fine-tuned (ft) models in the train-test match setting for the BBN, LDC, and NormLDC encodings across all difficulty levels. Best performance for each encoding and difficulty level within o and ft models (separately) is bolded and second best, underlined, the differences of which are all statistically significant ($p<0.001$) except for ties.

### v4

@@ 4.1 Experiment 1: Baselines
Table 2: Bootstrapped test performance (AUC) across all out-of-the-box (o) models for the BBN, LDC, and NormLDC encodings across all difficulty levels. Best performance for each encoding and difficulty level is bolded and second best, underlined, the differences of which are all statistically significant ($p<0.002$) except for ties. The largest standard error is 0.0004.
@@
For the TF-IDF baseline, the vectorizer was fit to the Reuters-21578 corpus, which contains 10,788 news documents and totals 1.3 million words (Lewis, 1997), [^10] We also tried fitting the vectorizer to the training set transcripts but found worse results so fit to Reuters to focus on (better) out-of-the-box performance of written text models. the document-term matrix obtained for each transcript in the test set trials, and the cosine similarity calculated between each matrix in each trial. For the out-of-the-box PAN baseline, PANgramso was trained on the most recent openly available PAN authorship verification dataset, fanfiction stories (Bevendorff et al., 2020), and evaluated on the test set trials using cosine similarity.
@@
AUC score test set results are in Table 2 and EER results are in Table 8 in Appendix A. Focusing on the leftmost column (BBN ‘text-like’ encoding, ‘base’ difficulty), there is a clear ranking in performance: PANgramso achieves the highest performance followed by TF-IDFo, LUARo, SBERTo, and CISRo, with all scores well above chance (AUC = $0.5$). AdHominemo performs the worst but still above chance. These initial results—considering only the ‘base’ setting for now—suggest at least some transfer from the text domain to the speech domain. However, we revisit this idea in Section 4.3 when discussing results on topic-controlled datasets.
@@ 4.3 Experiment 3: Topic
To test the effect of topic, after running Experiment 1 on the ‘base’ dataset, we ran it on the ‘hard’ and ‘harder’ datasets. Since performance of both LUARo and LUARnormo was highest on the NormLDC encoding previously, we included it here as well. The full AUC results across all settings are shown in Table 2 and the corresponding EER results are in Table 8.
@@ Appendix A Appendix
Table 8: Bootstrapped test performance (EER) across all out-of-the-box (o) models and fine-tuned (ft) models in the train-test match setting for the BBN, LDC, and NormLDC encodings across all difficulty levels. Best performance for each encoding and difficulty level within o and ft models (separately) is bolded and second best, underlined, the differences of which are all statistically significant ($p<0.002$) except for ties.